# Natural Language Processing RELIES on Linguistics

### v3

@@ 3 Resources
According to some predictions, half of the world’s spoken languages might be on the brink of extinction (UNESCO, 2023), prompting the development of computational tools to document and revitalize endangered languages (e.g., van Esch et al., 2019; Levow et al., 2021; San et al., 2022). The lack of existing language resources in endangered languages can also frustrate the development of NLP technologies that could serve communities with ties to such languages.
To counteract the ongoing depletion of the world’s languages, collaborations involving community members, linguists, and NLP practitioners seem particularly promising. Linguistic expertise is relevant for structurally collecting and documenting language data, as well as helping to distill and compress knowledge about the language into human dictionaries and rules. This linguistic knowledge may also be relevant then when building parallel corpora from which we might even then create machine translation methods to compress language knowledge into a neural network’s parameters, as hypothesized by Bird and Chiang (2012); Abney and Bird (2010).
Such projects are by nature highly interdisciplinary, as NLP experts and linguists rely on each others’ input. Recent work has brought community members and NLP practitioners together for the purpose of language documentation through provisioning access to speech tools such as ASR (San et al., 2022), building datasets on online platforms (Zuckermann et al., 2021; Everson et al., 2019; Dunham et al., 2014), analysis of phonetic data (Kempton, 2017), and linguistic feature structures to describe “particularly morphologically complex, low-resource languages” (Van Gysel et al., 2021). As endangered languages tend to be under-resourced, we return to this topic in §5.
Resources with linguistic annotations
These encompass datasets of various sorts of grammatical and semantic structures. Linguistically detailed resources prominent in current NLP research include Universal Dependencies treebanks, with morphosyntactic annotations for over a hundred languages (UD; Nivre et al., 2016), and Abstract Meaning Representation, which is a framework for describing semantic graphs that now underlies datasets in several languages (AMR; Banarescu et al., 2013; Wein and Schneider, 2024). Linguistic knowledge and awareness is, of course, needed for designing these frameworks, but also for conducting annotation and verifying model output. UD and AMR are applied linguistic theories (in the sense of Raskin, 1985, p. 275).
Research in and with such frameworks continues apace, with numerous recent workshops in the ACL community focused on grammar and semantic frameworks (Bonial and Tayyar Madabushi, 2023; Dakota et al., 2023; Bonn and Xue, 2023; Grobol and Tyers, 2023; Rambow and Lareau, 2023) and more than 100 publications addressing applications of the AMR framework alone [^7] https://nert-nlp.github.io/AMR-Bibliography/. Recent areas of NLP where integration of AMR has seen success include, i.a., machine translation (Song et al., 2019; Nguyen et al., 2021; Li and Flanigan, 2022), human-robot interaction (Bonial et al., 2019; Abrams et al., 2020; Bonial et al., 2020, 2021, 2023), and information extraction (Garg et al., 2016; Wang et al., 2017; Rao et al., 2017; Zhang et al., 2021; Yi et al., 2024). For a comprehensive overview of AMR applications, we refer the reader to two recent surveys: Wein and Opitz (2024) and Sadeddine et al. (2024).
Still, though interest in structured linguistic representations does not seem to be receding, it is not clear yet which avenues will be most fruitful for downstream application of formal representations. While some approaches have successfully leveraged explicit linguistic structures alongside neural models for applications (e.g., Kapanipathi et al., 2021; Li and Flanigan, 2022), neuro-symbolic approaches do not consistently outperform strictly neural approaches (Hamilton et al., 2022; Shwartz, 2023). [^8] Still, as two interesting recent developments regarding neuro-symbolic models, we see the creation of challenging datasets that reveal the need for their application (Fodor et al., 2024), and abstract symbolic cross-lingual representations that even enable fully-symbolic cross-lingual comparison (Zhang et al., 2024b). We consider the utility of such formal representations for evaluating and interpreting models (e.g., Xu et al., 2021; Opitz and Frank, 2022) to be an exciting path. Another exciting application lies in using NLP models for studying questions about language, which we will continue to discuss below.
4 Evaluation
We illustrate where linguistic influence is useful to successfully evaluate an NLP system, that is, to describe the degree to which a system conforms to our expectations of accuracy on a task.
A common practice in NLP system evaluation is to apply one or several automatic metrics that assess model output against a reference, ideally supplemented by a human evaluation. As we discussed in §3, expertise in computational linguistics can be important to designing crowdsourcing protocols which lead to collection of accurate reference annotations. Similarly, human evaluation studies require language knowledge that ensure effective error analysis and quality assessment (Freitag et al., 2021a).
Gold standard evaluation
Linguistic resources, particularly so-called gold standard annotations, play an important role in the evaluation of NLP technologies. For instance, they contribute to the comprehensive benchmarking of general AI systems Tenney et al. (2019b); Pimentel et al. (2020a), where many of the tasks packaged in the Beyond the Imitation Game Benchmark (BIG-bench), a large test-suite for tracing AI development Srivastava et al. (2023), are linguistic tasks, such as bridging resolution and detecting common morphemes. In low-resource and extinct languages, models are benchmarked with linguistic puzzles (Bean et al., 2024); this draws a connection to another area where NLP relies on linguistics (low-resource settings), which we will cover later in more detail (reLies, section 5).
Linguistic resources are also crucial for system diagnostics which is a part of evaluation, but does not necessarily have the aim to achieve a “global ranking” (benchmarking) of AI systems. Instead, diagnostic methods help find more targeted perspectives for system improvement, with possible relevance even to broader society, since linguistic phenomena are often contingent with issues of societal interest, as is the case of coreference and gender bias (Rudinger et al., 2018) or question answering with social biases Parrish et al. (2022).
Overall, linguistic diagnostics can come in many forms, and here we give a non-comprehensive overview. E.g., with their ‘CheckList’ approach, Ribeiro et al. (2020) test NLP systems in various categories, leveraging linguistic features (negation, part of speech, sentiment, etc.) to build templates. In a similar style, Song et al. (2022) create a data set for assessing LLM performance on various Chinese linguistic phenomena, and Parcalabescu et al. (2022) employ linguistic knowledge to design image-caption foils that reveal the behavior and understanding of V&L systems in different visual-linguistic aspects. Moore (2009) outlines examples of challenges for models which require linguistic knowledge to unpack, such as structural parallelism and translation of “WH” questions.
Automatic diagnostics for text generation systems can be conducted through controlled and targeted measurements on parsed semantic role Lo and Wu (2011); Fan et al. (2023) or semantic (sub-)graph structures Birch et al. (2016); Opitz and Frank (2021); Opitz (2023) (which calls for research to improve the intrinsic NLP task of parsing). Finally, there is the large area of probing systems with linguistic structures Tenney et al. (2019c); Rama et al. (2020); Pimentel et al. (2020b); Starace et al. (2023), possibly tailored at certain linguistic branches, e.g., psycholinguistics Gauthier et al. (2020).
In sum, linguistics helps to take a system’s fingerprint, evaluate a system in particular categories, fosters understanding of complex models by binding observed behavior to interpretable linguistic categories (cf. Interpretability, §6), and helps to build resources (cf. Resources, §3).
Human evaluation
Despite being an important component of model evaluation, human evaluation practices are currently varied. The combination of varied human evaluation practices with progress towards highly fluent models has led to a degradation of human evaluations, meaning that not all human evaluation studies are useful measures of model output (Clark et al., 2021; Freitag et al., 2021a).
Whether making use of implicit linguistic knowledge held by native speakers (e.g. judging the fluency of system output) or more explicit linguistic training (e.g. assessing the syntactic diversity of system output), human evaluations serve a crucial role in reliably assessing the state of the field (Michael et al., 2016; Callison-Burch et al., 2008), especially as systems continue to improve (Ma et al., 2019); this therefore motivates the use of human evaluation and its standardization with regard to design and terminology (Howcroft et al., 2020; Belz et al., 2020; van der Lee et al., 2019).
Meta-evaluation
In addition to their use in human evaluation studies, human judgments are often used to train metrics as well as to judge the quality of automatic metrics themselves (“meta-evaluation”). Typically, this is done by correlating agreement between the human judgments and the automatic metric scores, a protocol that has been used for numerous automatic metrics, e.g., in the Machine Translation metric shared tasks Ma et al. (2019); Mathur et al. (2020), the WebNLG challenge Gardent et al. (2017); Zhou and Lampouras (2020), generation of medical notes Moramarco et al. (2022), and generation of meaning graphs Manning et al. (2020).
To design a meta-evaluation framework for metrics that assess generated language, expertise in linguistic theories may be required for some tasks, and possibly offer general value due to training in abstract thinking. For instance, similar to how training in mathematics would help in solving machine learning problems, training in linguistics can help study and diagnose complex behavior of models and metrics that produce and evaluate language data; this is because the specific linguistic phenomena which may be more challenging for the models need to be understood in order to be identified, such as in the case of probing BERT via psycholinguistic diagnostics (Ettinger, 2020). Further, in their WMT shared task overview paper on MT metrics, Freitag et al. (2021b) find that most metrics struggle to handle complex semantic phenomena as well as to appropriately punish the reversal of sentiment or negation; this is established through the use of test sets which assess specified linguistic features, such as negation and sentiment polarity.
One way to circumvent the need for meta-evaluation is with a ‘human user in the loop’ Bodén et al. (2021), though downstream evaluations may be costly and irreproducible in many circumstances.
New metrics
In some cases, linguistic information in the form of annotations can help build strong automatic metrics. For instance, entailment data [^9] Even though entailment annotations may come from non-linguists, the notion of textual entailment was inspired by traditions from logic and formal semantics Dagan et al. (2006); Bowman et al. (2015). is used to build accurate and efficient neural NLG evaluation metrics Chen and Eger (2023), including metrics that address the problem of checking whether a generated text is factual Xie et al. (2021); Scialom et al. (2021); Honovich et al. (2022); Steen et al. (2023). [^10] Factual text generation is a fundamental problem that may rise in importance against the backdrop of new risks through LLMs De Angelis et al. (2023); Gupta et al. (2023); Li et al. (2023), in particular with regard to the phenomenon of hallucination (Xu et al., 2024). For finer classification, pragmatics-informed annotations of veridicality could be crucial (de Marneffe et al., 2012).
Finally, we observe that measures for biases of other sorts, like social biases, have drawn upon the efforts of linguists. For example, Steen and Markert (2024) propose bias evaluation measures that leverage socio-linguistic lexicons (gender, race) and OntoNotes (Pradhan et al., 2013) for template creation. This work also highlights the potential use of linguistically annotated corpora in constructing evaluation data and evaluation measures.
5 Low-Resource Settings
The term “low-resource” can have multiple meanings Hedderich et al. (2021). In NLP and ML generally, an important problem is efficiently learning to generalize to new tasks, domains, and languages, where the term “new” is not to be understood in a strict way, but can in this context also indicate a situational lack of information, e.g., little, or no available training data or lack of feedback of some other sorts. Specifically, “new” languages would include endangered languages (as discussed in §3 above), and extinct languages, as well as languages that are widely used but not as commercially prioritized for language technologies Joshi et al. (2020); Petrov et al. (2023); Khanuja et al. (2023); Hu et al. (2023), i.e., all global and historic languages. In such settings, often we lack the direct feedback that we usually use to train or create NLP system in higher-resource settings. As an edge case, we’d have only archaeological fragments and perhaps dictionaries available.
Next, within the broader field of computer science, the term low-resource can mean that a method can be run with strictly limited compute-resources. These methods would, e.g., offer quicker inference, or have a smaller energy footprint.
Finally, we touch on the need for linguistically sensitive supervision when it comes to developing language technologies for under-resourced languages, as it concerns the native speakers of the language.
Processing of global and historic languages
Linguistics is crucial in situations where the amount of data is limited, e.g., when developing technologies for languages with a limited amount of available recordings or written data. This could include languages that lack digitized text (and/or even) audio/video recordings, or have only a few and noisy transcriptions of recordings Bird (2020). The only few bits of information readily at our hands could consist of, e.g., archaeological fragments or descriptions of the language and contexts, for instance in the form of dictionaries written by (field-)linguists. In these and similar scenarios, using linguistic input can improve generalization performance across an array of LLM models, as demonstrated by Zhang et al. (2024a) who exploit linguistic features and Lu et al. (2024) who use dictionaries for balanced data augmentation; these findings also generalize to audio (Kim et al., 2024).
What may appear first as a potential ‘linguistics-free’ avenue for addressing some of these extremely low resource-scenarios, is Tanzer et al.’s (2024) study that seems to show that LLMs can to a certain extent translate an under-resourced language from only one linguist’s grammar book, seemingly calling into question the assumption that “language models would have to wait for a usable language corpus to be created” (Lewis, 2023). Still, this work relies on linguistics in several crucial aspects: a linguist is required to reasonably assess LLM performance (here, to establish an upper bound), and even the largest LLMs considerably underperform relative to a linguist that has read the book but was unfamiliar with the language. Of course, it is also the case that this fundamental resource was created by a linguist.
Processing particular types of language/particular languages can also pose challenges that require input from linguistics. For examples, polysynthetic languages such as those spoken in Canada (Gupta and Boulianne, 2020a, b) pose problems for automatic speech recognition (ASR): if ‘word’ is taken as the basic unit for recognition (which makes sense for most languages, but not these ones) the out-of-vocabulary (OOV) rate on new data is extreme. As a result, ASR researchers need linguistic knowledge in order to effectively design their systems to account for the diversity of global languages—e.g. by breaking long polysyllabic words up into morphemes or syllables. Note that this isn’t the only problem for NLP with indigenous polysynthetic languages—many are also under-resourced. Finite-state methods can be superior to machine learning methods for many polysynthetic and other morphologically rich languages, at least with respect to their precision (Forbes et al., 2021). They can also provide the groundwork for leveraging neural advances by generating silver annotations (Rueter and Hämäläinen, 2020; Rueter et al., 2023).
Furthermore, even in the case where there is much (but not abundant data) available, linguistics can be of value: It can inform data selection in multi-lingual pre-training Ogunremi et al. (2023), or enable more language-balanced tokenization with morphemes Creutz and Lagus (2002); Limisiewicz et al. (2024). While it may be possible to pretrain multilingual LLMs on data from hundreds of languages (e.g., Imani et al., 2023), experiments on unseen languages indicate that “simply scaling up the number of languages […] in the pre-training is unhelpful” for out-of-sample generalization (Adelani et al., 2024), and “bigger is not always better” Wilcox et al. (2024), suggesting the models may benefit from incorporation of a linguistic signal. It may also be possible to leverage auxiliary linguistic structures to improve upon a pretrained language model (Prange et al., 2022).
@@ 5 Low-Resource Settings
Systems that require less compute power, particularly at inference time, are also of increasing interest. Critically, off-the-shelf conventional parsers/taggers clearly defend their status as the go-to method in targeted situations, in academia and industry. They are ideal for conducting a lightweight data analysis or fast (and affordable) processing of huge corpora, e.g., up until extracting relation networks with linguistic machinery Angeli et al. (2015); Martinez-Rodriguez et al. (2018), justifying future work into improving the breadth and quality of the computational linguistic tool box.
More generally, there are several scenarios where we might want to reduce the time of inference Treviso et al. (2023) or the size/energy footprint of a model Strubell et al. (2019). For many end-user applications, sending data to an LLM (or hosting an LLM locally) may also simply not be practical, due to issues of connectivity, privacy, and runtime.
Some effective approaches to computational efficiency are pure engineering solutions, such as quantization Dettmers et al. (2022) or low-rank adaptation Hu et al. (2022) of LLMs. However, soft inductive linguistic biases might be complementary, or an alternative in specific situations (we refer back to the paragraph above for examples). To find a right balance of linguistic biases and pure LLM engineering, new evaluation metrics might be useful that facilitate such investigations into resource-performance tradeoffs Zhou et al. (2022).
@@
Developers of technologies for under-resourced languages (e.g., a QA dialog system for a language with only a handful of speakers) are bound to not only encounter data scarcity, but also scarcity of input from native speakers who could help oversee a system’s design and application. An outcome that should be prevented is triggering ethical issues in local communities that can arise when technology is released in a top-down fashion Bird (2022); Doğruöz and Sitaram (2022).
To reduce the risk of harmful biases or misunderstandings, and when applied with care, linguistic sensitivity through pragmatics and field linguistics can help to better become aware of the diverse cultural contexts of local language communication situations. Thus, while in §3 we outlined the importance of (interdisciplinary) linguistics for collecting and documenting data to preserve under-resourced languages, here we argue that developing and applying text technologies for under-resourced languages also profits from linguistic input.
@@ 6 Interpretability and Explanation
How can we efficiently understand, reason about, and explain observations of observed language phenomena, language systems, and processes of language? For this, it is necessary to have a shared terminology and metalanguage, onto which we can bind our observations, and develop binding methods that relate model processes and internals to human-understandable concepts. We argue that linguistics provides valuable tools for helping with both aspects.
We find that NLP is already permeated by a linguistic metalanguage. This metalanguage is employed by researchers, laypeople, and users to talk and reason about their observations, such as model predictions, and hypothesizing explanations. As a simple example, take the grammatical concept of noun: even if we all may not agree in every instance on what a noun exactly is, the approximations that we have of this concept constitute a common ground that can be useful, e.g., to describe observed model behavior (“the model is mostly attending to nouns for its predictions”). While noun is a concept familiar to anybody who has used a dictionary, exposure to linguistics enriches one’s lexicon with terms like syntactic ambiguity, presupposition, code-switching, and polysynthetic language. These terms of art reflect a level of understanding of language systems—how they are organized, how they vary, and so on.
@@
Fundamentally, with a precise metalanguage for the inner workings of language, it is possible to formulate more rigorous hypotheses and claims about language models. This can be seen in the literature in observations like “phonemic representations exhibit higher similarities between languages compared to orthographic representations” (Jung et al., 2024) and “difficult to score [in MT evaluation] are the transitive past progressive, multiple connectors, and the ditransitive simple future I for English to German, and pseudogapping, contact clauses, and cleft sentences for English to Russian” (Avramidis et al., 2024).
@@
For understanding computational models of language such as LLMs with billions and trillions of parameters, directed binding methods are being developed that learn to relate model internals and processes to specific human-understandable descriptions, that often lie within our vocabulary of linguistic meta-language. We will consider a few examples of such methods that particularly use a linguistic notion of binding: Rassin et al. (2023) bind linguistic concepts in V&L diffusion models leveraging syntax trees and attention maps; Haviv et al. (2023) show how to understand transformer memories with idioms, or detecting linguistic feature representations in LLMs with prompts Di Marco et al. (2023); Jumelet and Zuidema (2023); Opitz and Frank (2022) enforce binding in a transformer model by relating parts of neural representations to semantic features, for greater interpretability of decisions. Explanations and binding can also adopt an outside, behavioral view Beguš et al. (2023); Behzad et al. (2023); Chang and Bergen (2024). E.g., Muñoz-Ortiz et al. (2023) contrast linguistic patterns in human- and LLM-generated text in morphological, syntactic, and sociolinguistic aspects. Since such methods can be also seen as diagnostic methods, the section on Evaluation (§4) is also relevant.
@@
Indeed linguistics, and binding observed behavior to said meta-language, can help us understand, or establish hypotheses, for what happens in complex NLP systems, and how they differ. A popular example is the hypothesis that a text Transformer model can be viewed through the lens of a classical NLP pipeline, which has received arguments for and against (Tenney et al., 2019a; de Vries et al., 2020; Niu et al., 2022). Alternatively, neural decisions can be viewed as constructed from interpretable concepts in a vocabulary Geva et al. (2022), or the learning trajectory of LLMs can be examined through a linguistic lens, such as noting the point when a model acquires syntax (Chen et al., 2024).
@@
Tying interpretability more immediately back to Resources (§3), we can use data sets that elicit linguistic phenomena to “open the black box” of LLMs down to the level of single neurons, as exemplified by Niu et al. (2024), who study neuron-level knowledge with the BLiMP corpus of linguistic minimal pairs (Warstadt et al., 2020). Through designing a resource using linguistically informed counterfactuals, Wang et al. (2024) show that BERT-like models primarily focus on sentence-level features, whereas LLMs such as GPTand Llama are sensitive to conventions, accuracy, language complexity, and organization. Experiments with complex and simple language structures can illuminate the settings in which LLMs learn successfully (Qin et al., 2024).
@@
Linguistically-agnostic interpretability approaches can be mechanics-based, i.e., formalizing the algorithms learned by neural networks Weiss et al. (2021); or eliciting attributions to input features, e.g., by employing integrated gradients Sundararajan et al. (2017), classifiers Ribeiro et al. (2016), or Shapley values Shapley (1951). However, still, for interpreting and evaluating the results of such methods, we depend on linguistic analysis Schuff et al. (2022); Feldhus et al. (2023); Moeller et al. (2024), to verify that a computational interpretability method “outputs meaningful explanations” Mardaoui and Garreau (2021) or “emphasize[s] specific types of linguistic compositions” (Kobayashi et al., 2024). This means that linguistic analyses can be used to validate interpretability methods whose algorithms do not explicitly incorporate such representations.
@@
Finally, at the thousand-foot view, we see ideas from linguistics and adjacent fields take center stage in debates about how to interpret what NLP models are capable of representing—like how to define machine ‘understanding’ (Dunietz et al., 2020; Ray Choudhury et al., 2022) and whether grounding is required for a model to capture meaning (e.g., Bender and Koller, 2020; Merrill et al., 2021; Pavlick, 2023).

### v4

@@ 5 Low-Resource Settings
According to some predictions, half of the world’s spoken languages might be on the brink of extinction (UNESCO, 2023). Thus, there is scope for NLP tooling to document and revitalize endangered [^9] Terms like ‘endangered’ and ‘documentation’ are widespread in the literature we are surveying, though we also wish to point at a critique of this terminology, suggesting ‘reclamation’ instead (Leonard, 2020). languages (e.g., van Esch, Foley, and San, 2019; Levow, Ahn, and Bender, 2021; San et al., 2022), ideally through collaborations involving community members, linguists, and NLP practitioners. Linguistic expertise is relevant for structurally collecting and documenting language data, as well as helping to distill knowledge about the language into dictionaries and rules. Recent work has brought community members and NLP practitioners together for the purpose of language documentation through provisioning access to speech tools such as ASR (San et al., 2022), building datasets on online platforms (Everson, Honore, and Grimm, 2019; Zuckermann et al., 2021), and analysis of phonetic data (Kempton, 2017).
@@
Systems that require less compute, particularly at inference time, are of great interest. Critically, off-the-shelf conventional parsers/taggers defend their status as the go-to method in some situations, like for a certain forms of data analysis (e.g., counting the most common nouns), or affordable processing of huge corpora (e.g., from sentence segmentation to OpenIE (Angeli, Johnson Premkumar, and Manning, 2015)).
Effective engineering solutions to efficiency are achieved through quantization Dettmers et al. (2022) or low-rank adaptation Hu et al. (2022) of LLMs. However, soft inductive linguistic biases might be complementary, or an alternative in specific situations. Tailored evaluation metrics could help assess any incurred tradeoffs Zhou et al. (2022).
@@
Developers of technologies for under-resourced languages do not only encounter data scarcity, but probably also scarcity of input from native speakers who could help oversee a system’s design and application. Indeed, releasing technology in a top-down fashion can be harmful to local communities Bird (2022); Doğruöz and Sitaram (2022); Bird and Yibarbuk (2024). Thus, when applied with care, (field-)linguistics can help become aware of the diverse cultural contexts of local language communication situations, resulting in more sensitive NLP technologies, or the abstention from releasing harmful ones (e.g., when there is no community oversight/feedback).
@@ 6 Interpretability and Explanation
How can we efficiently reason about, and explain observations of observed language phenomena, language systems and processes, as well as any of their models in the form of NLP systems? We need a shared terminology and metalanguage, and develop binding methods that relate model internals to human-understandable concepts from said terminology. We argue that linguistics fundamentally helps with both aspects.
@@
As the field concerned with systems of human language, linguistics offers a metalanguage—in the form of jargon and formalisms—for precise characterization of NL phenomena. The field of NLP draws on this metalanguage, as can be seen in the literature in observations like “phonemic representations exhibit higher similarities between languages compared to orthographic representations” (Jung et al., 2024) and “difficult to score [in MT evaluation] are the transitive past progressive, multiple connectors, and the ditransitive simple future I for English to German, and pseudogapping, contact clauses, and cleft sentences for English to Russian” (Avramidis et al., 2024).
@@
For understanding computational models of language such as LLMs with billions and trillions of parameters, directed binding methods are being developed that learn to relate model internals and processes to specific human-understandable descriptions, that often lie within our vocabulary of linguistic metalanguage. We will consider a few examples of such methods that particularly use a linguistic notion of binding: Rassin et al. (2023) bind linguistic concepts in V&L diffusion models with syntax trees and attention maps; Di Marco, Hämmerl, and Fraser (2023) and Jumelet and Zuidema (2023) detect linguistic feature representations in LLMs with prompts; Opitz and Frank (2022) bind parts of neural embedding representations to semantic aspects; Geva et al. (2022) view model decisions as constructed from concepts in a vocabulary. Explanations and binding can also adopt an outside, behavioral view Beguš, Dąbkowski, and Rhodes (2023); Behzad et al. (2023); Chang and Bergen (2024). For instance, Muñoz-Ortiz, Gómez-Rodríguez, and Vilares (2023) contrast linguistic patterns in human- and LLM-generated text in morphological, syntactic, and sociolinguistic aspects. Since such methods can be also seen as diagnostic methods, the section on Evaluation (section 4) is also relevant.
@@
On a bigger picture, linguistic metalanguage can help us understand, or establish hypotheses, for what happens in complex NLP systems, and how they differ. An example is the hypothesis that a pre-trained LM can be viewed through the lens of a classical NLP pipeline, which has received arguments for and against (Tenney, Das, and Pavlick, 2019; de Vries, van Cranenburgh, and Nissim, 2020; Niu, Lu, and Penn, 2022). The learning trajectory of LLMs can be traced linguistically, e.g., noting the point of syntax acquisition (Chen et al., 2024). LLMs have also been interpreted as “models of varieties of language” (a sociolinguistic notion, Grieve et al., 2025).
@@
Building a bridge to Resources (§3), we can use data sets that elicit linguistic phenomena to “open the black box” of LLMs down to the level of single neurons, as exemplified by Niu et al. (2024), who study neuron-level knowledge with the BLiMP corpus of linguistic minimal pairs (Warstadt et al., 2020). With similar means, Wang, Hu, and Zhao (2024) show that BERT-like models focus on sentence-level features, whereas LLMs such as GPT or Llama are sensitive to conventions, language complexity, and organization. Experiments with complex and simple language structures can illuminate learning environments for LLMs (Qin, Saphra, and Alvarez-Melis, 2024).
@@
Linguistically-agnostic interpretability methods can be mechanistic, i.e., formalizing the algorithms learned by neural networks Weiss, Goldberg, and Yahav (2021), or based on attributing a prediction to inputs via integration Sundararajan, Taly, and Yan (2017), classifiers Ribeiro, Singh, and Guestrin (2016), or Shapley values Shapley (1951). However, still, for interpreting and evaluating the results of such methods, we rely on linguistic analysis Schuff et al. (2022); Feldhus et al. (2023); Moeller, Nikolaev, and Padó (2024), to verify that an interpretability method “outputs meaningful explanations” Mardaoui and Garreau (2021) or “emphasize[s] specific types of linguistic compositions” (Kobayashi et al., 2024). This means that linguistic analyses are used to validate interpretability methods whose algorithms do not explicitly incorporate such representations.
@@
Finally, at the thousand-foot view, we see ideas from linguistics and adjacent fields take center stage in debates about how to interpret what NLP models are capable of representing—like how to define machine ‘understanding’ (Dunietz et al., 2020; Ray Choudhury, Rogers, and Augenstein, 2022) and if grounding is required for a model to capture meaning (e.g., Bender and Koller, 2020; Merrill et al., 2021; Pavlick, 2023).